Consider a binomial probability distribution with p 0.55 and n 7. Determine the probabilities below. a) P(x 2) b) P(xs1) c) Px>5) a) P(x = 2 (Round to four decimal places as needed.) b) Ps1)- (Round to four decimal places as needed.) c) P(X> 5)= □ (Round to four decimal places as needed.) Enter your answer in each of the answer boxes.
Answer #1

P(X) = nCx px qn-x

n = 7

p = 0.55

q = 1 - p = 0.45

a) P(X = 2) = 7C2 x 0.552 x 0.455

= 0.1172

b) P(X \leq 1) = P(0) + P(1)

= 0.457 + 7x0.55x0.456

= 0.0037 + 0.0320

= 0.0357

c) P(X > 5) = P(6) + P(7)

= 7x0.556x0.45 + 0.557

= 0.0872 + 0.0152

= 0.1024